OpenStudy (kawii2004):

Find the perimeter of a rectangle with l = 1.2 m and w = 0.6 m. A. 0.72 m B. 1.8 m C. 3.6 m D. 7.2 m

OpenStudy (kawii2004):

help

OpenStudy (anonymous):

Okay, here u go P=2(l+w)=2·(1.2+0.6)≈3.6m

OpenStudy (anonymous):

The perimeter is 2(l*w)

OpenStudy (kawii2004):

so whats the awnser

OpenStudy (kawii2004):

hello

OpenStudy (kawii2004):

hello i wanted to know

OpenStudy (anonymous):

Wait i will give u the answer

OpenStudy (kawii2004):

ok

OpenStudy (anonymous):

It is C) 3.6m

OpenStudy (kawii2004):

thank you so much

OpenStudy (anonymous):

Your welcome and if you need any help just call me

OpenStudy (anonymous):

:)

OpenStudy (kawii2004):

i love you for this i got a 100

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

U just made my day